# SS4Rec: Continuous-Time Sequential Recommendation with State Space Models

### v1

@@ Abstract.
Sequential recommendation is a key area in the field of recommendation systems aiming to model user interest based on historical interaction sequences with irregular intervals. While previous recurrent neural network-based and attention-based approaches have achieved significant results, they have limitations in capturing system continuity due to the discrete characteristics. In the context of continuous-time modeling, state space model (SSM) offers a potential solution, as it can effectively capture the dynamic evolution of user interest over time. However, existing SSM-based approaches ignore the impact of irregular time intervals within historical user interactions, making it difficult to model complexed user-item transitions in sequences. To address this issue, we propose a hybrid SSM-based model called SS4Rec for continuous-time sequential recommendation. SS4Rec integrates a time-aware SSM to handle irregular time intervals and a relation-aware SSM to model contextual dependencies, enabling it to infer user interest from both temporal and sequential perspectives. In the training process, the time-aware SSM and the relation-aware SSM are discretized by variable stepsizes according to user interaction time intervals and input data, respectively. This helps capture the continuous dependency from irregular time intervals and provides time-specific personalized recommendations. Experimental studies on five benchmark datasets demonstrate the superiority and effectiveness of SS4Rec.
@@
State Space Model, Continuous-Time Modeling, Sequential Recommendation
isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
@@
Recommendation systems are widely employed in social media and e-commerce to alleviate information overload (Wu et al., 2023). The core of recommendation systems is accurately predicting user interests in order to suggest appropriate items for users from a large candidate pool (Wang et al., 2019). Notably, user interests are always changing continuously over time, and this evolution is reflected in their historical interaction sequence. As a result, sequential recommendation systems have been developed to offer personalized online services by capturing the users’ evolving interests through their interaction sequences.
@@
Earlier sequential recommendation approaches adopt Markov chains to model sequential patterns by learning transition probabilities between items (Rendle et al., 2010; He and McAuley, 2016). With the rise of deep neural networks, a series of related methods emerged, including those based on recurrent neural networks (RNNs) (Jain and Medsker, 1999; Hidasi et al., 2016; Tan et al., 2016), convolutional neural networks (Tang and Wang, 2018), and Transformer models (Vaswani et al., 2017; Kang and McAuley, 2018; Sun et al., 2019). Although these approaches achieve promising performance, they are unable to effectively capture long-range dependencies and suffer from the inference inefficiency problem. Additionally, these models are inherently discrete and designed for dealing with discrete systems, however, user interests evolve continuously and vary over time, which requires recommendation systems to adapt to the user’s current state as well as the dynamic interest evolution.
@@
Recently, state space models (SSMs) have become an alternative to neural network-based models for their ability to model continuous dynamic systems. S4 (Gu et al., 2022) and S5 (Smith et al., 2023) utilize HIPPO matrix initialization and parallel scanning for long-range reasoning and fast inferring, respectively. The latest SSM, Mamba (i.e., S6) (Gu and Dao, 2024), leverages the input-dependent selective state space model to memorize input-relevant information and filter out irrelevant information. This enhances its sequential modeling capability and encourages researchers to explore the potential of SSMs for sequential recommendation. Mamba4Rec (Liu et al., 2024) and RecMamba (Yang et al., 2024) employ Mamba to capture sequential dependencies from historical interactions to model complex relationships between users and items and address the dilemma of recommendation performance and inference efficiency to a certain degree.
@@
Despite the success of SSMs in efficient sequential recommendations for long sequences or large-scale datasets, these methods are still under the assumption of uniform time intervals since Mamba was initially designed for NLP tasks. Therefore, existing SSM-based approaches focus on capturing relation-aware information but they neglect the effect of varying temporal information contained in user interest evolution. In reality, the observed interactions are often recorded as discrete time series data with irregular time intervals (Chen et al., 2024), which are crucial for maintaining the underlying continuity. Even though some methods (Li et al., 2020; Ye et al., 2020; Zhang et al., 2023; Wu et al., 2020) introduce time-dependent embeddings or kernel functions to model temporal information, they still struggle to overcome the assumption of uniform time intervals. This assumption will damage the underlying continuity of the recommender, hindering its ability to capture users’ dynamic evolution interests and their time-specific preferences (e.g., morning and evening).
@@
To address the issues above, we observe that different SSM architectures own distinct strengths and propose a hybrid SSM-based model SS4Rec for continuous-time sequential recommendation. SS4Rec consists of a time-aware SSM and a relation-aware SSM. The discretized steps of the two SSMs are based on the time interval sequences and the user interaction sequences, respectively, in order to capture the user interests from different aspects. Specifically, SS4Rec first leverages the time-aware SSM as an effective encoder to handle variable observation intervals, capturing time patterns. Then, it incorporates the selective mechanism into the relation-aware SSM to learn important and relevant information, thereby further managing sequence dependencies. In addition, the residual connection and layer normalization techniques are employed to prevent network degradation and maintain stability during training.
Major contributions of this work are: (1) The sequential recommendation problem is formulated as a continuous time-varying system and is solved by state space models with time-varying parameters. This work breaks the assumption of uniform time interval in previous works, enabling to model user interest evolution with irregular-interval interactions directly, providing more accurate and time-specific recommendations. (2) SS4Rec is the first work that explores the potential of integrating distinct parameterized state space models to capture the continuity of the recommendation system from both temporal and sequential perspectives simultaneously, meanwhile exerting the advantages of SSMs on long-range dependencies and inference efficiency. (3) Extensive experiments are conducted on five real-world datasets, and the results show the superiority and effectiveness of SS4Rec compared with other state-of-the-art models.
2. Related Work
2.1. Sequential Recommendation
@@
A variety of sequential recommendation methods have been developed over time. Early works on sequential recommendation systems typically adopted the Markov chain, for example, FPMC (Factorizing Personalized Markov Chains) (Rendle et al., 2010) combines the factorization machine and Markov chain techniques to calculate a personalized item-item relationship transition matrix for each user. This enables the prediction of a user’s next interaction based on the last interaction.
@@
In recent years, deep neural networks (DNNs) have garnered much attention due to their powerful expressive capabilities. Some DNN-based models have emerged in the field of sequential recommendation. GRU4Rec (Hidasi et al., 2016) utilizes Gate Recurrent Unit (GRU) (Cho et al., 2014) network with ranking loss to capture the user interests within the sequence. NARM (Li et al., 2017) applies GRU to extract the users’ general interests and their intentions from the global encoder and local encoder respectively. SASRec (Kang and McAuley, 2018) employs the Transformer structure to model user interaction sequences which captures long-term dependencies and adaptively adjusts the attention weights of items in the sequence, thereby focusing on the important information. Bert4Rec (Sun et al., 2019) introduces a Cloze task and utilizes a bidirectional attention mechanism to learn contextual information in the interaction sequences. DCRec (Yang et al., 2023) and GCL4SR (Zhang et al., 2022) leverage self-supervised learning techniques to capture complex relationships within sequential data and improve recommendation performance. All of the above mentioned approaches devote to modeling sequential dependencies among historical user behaviors but ignore the timestamps of interactions impeding their ability to capture the dynamic evolution of user interests.
2.2. Time-Aware Recommendation
@@
There exist several time-aware sequential recommendation approaches that consider temporal information in the learning of user preferences. TiSASRec (Li et al., 2020) adopts a time-aware self-attentive model to learn the weights of different items in a sequence, integrating the absolute position and the relative time interval between items to predict future items. TGSRec (Fan et al., 2021) unifies sequential patterns and temporal collaborative signals by introducing a Temporal Collaborative Transformer (TCT) layer, capturing both user-item interactions and temporal dynamics in continuous-time recommendation. TAT4SRec (Zhang et al., 2023) processes item sequences and timestamp sequences separately and utilizes a window function to preserve the continuity in timestamps. TASER (Ye et al., 2020) introduces a relative temporal attention net to capture the pairwise dependencies using absolute temporal vectors. The work of TiCoSeRec (Dang et al., 2023) verifies that uniformly distributed time intervals enhance performance. Different from these time-discretized approaches, we explore the continuous-time recommendation system and try to mine dynamic evolving information underlying continuous systems.
2.3. State Space Model
@@
State space models have emerged as a promising class of architectures, which achieve excellent performance in sequence modeling tasks (Xu et al., 2024). The first structured state space sequence model S4 (Gu et al., 2022) addresses the issue of capturing long-range dependencies with the state matrix initialization by HiPPO framework (Gu et al., 2020). Different from the single-input, single-output scheme of S4, S5 (Smith et al., 2023) adopts multi-input, multi-output (MIMO) SSM. Based on a diagonal state matrix, S5 has linear computational complexity and time-varying parameterization via parallel scans. The latest work of SSM, Mamba (Gu and Dao, 2024), integrates with a selection mechanism to filter irrelevant information and memorize relevant information. It also designs a hardware-aware algorithm for both efficient training and inference. Mamba2 (Dao and Gu, 2024) connects attention variants with selective SSM directly for large-scale language modeling. Jamba (Lieber et al., 2024) is a hybrid architecture for large-scale modeling that combines Mamba with Transformer and mixture-of-experts module, to take full advantage of each architecture.
@@
Recently, a series of works introduce Mamba into recommendation systems owing to its impressive performance. Mamba4Rec (Liu et al., 2024) is the first work to employ the Mamba block for sequential recommendation with improving performance and fast inference. RecMamba (Yang et al., 2024) utilizes the Mamba in place of the Transformer block in SASRec for lifelong sequences. MaTrRec (Zhang et al., 2024) combines Mamba with Transformer to incorporate their strengths of capturing long-range and short-range dependencies. SSD4Rec (Qu et al., 2025) introduces bi-directional SSM block to capture contextual information, likes Bert4Rec. Time-specific personalized recommendations with irregular intervals are not fully explored by previous research efforts. By formulating this problem as a continuous time-varying system and adopting hybrid state space models to capture both temporal and sequential information of user interactions, our work sheds light on the continuous-time sequential recommendation.
3. SS4Rec
3.1. Framework Overview
@@
In recommendations, the dynamic changes of user interests can be treated as a continuous time-varying system, user-item interaction sequences depict this system from discrete states with irregular time intervals. Our proposed SS4Rec concurrently models the relationship between user-item interactions (i.e., taking as input points of system) and captures the dynamic evolution of the continuous-time system by designed time-aware SSM and relation-aware SSM.
@@
An overview of the SS4Rec framework is shown in Figure 1. SS4Rec is a hybrid model based on two different types of SSM for sequential recommendation. It involves initialization, SS4Rec blocks, and continuous-time prediction. Specifically, each SS4Rec block comprises two principal components: (1) Time-Aware SSM discretized by variable time steps; (2) Relation-Aware SSM with a selection mechanism. Residual connection and layer normalization operations are introduced between each layer. Next, we will present the details of our model.
[…]
Figure 1. Framework of SS4Rec.
3.2. Initialization
@@
In the recommendation system, $\mathbf{S}_{u}=[v_{1},v_{2},\dots,v_{n}]$ represents the item sequence of user $u$ and $\mathbf{T}_{u}=[t_{1},t_{2},\dots,t_{n}]$ represents the associated time sequence of user $u$, where $v_{i}$ denotes the ID of $i$-th item and $t_{i}$ denotes the interaction timestamp of $v_{i}$.
@@
In the beginning, the user interaction information is initialized from two aspects. One is the initialization of the item sequence, which is mapped into $D$-dimensional embedding vectors using an embedding layer. In the embedding layer, a learnable item embedding matrix $\mathbf{E}\in\mathbb{R}^{N\times D}$ is employed to project item IDs to high-dimensional vectors, where $N$ refers to the number of whole items and $D$ denotes the dimension of item embedding vectors. The embedding layer transforms each sequence $\mathbf{S}_{u}$ to item embedding vectors $\mathbf{X}_{u}\in\mathbb{R}^{n\times D}$. The other is the initialization of the time interval sequence. According to the time sequence $T_{u}$, we can compute the time intervals $\Delta_{i}=|t_{i+1}-t_{i}|$ between two items $v_{i}$ and $v_{i+1}$. Hence, the time interval sequence of user $u$ can be obtained as $\mathbf{\Delta T}_{u}=[\Delta_{1},\Delta_{2},\dots,\Delta_{n}]\in\mathbb{R}^{n}$.
@@
For efficient computing, we process sequences in a batch to generate output embeddings $\mathbf{X}\in\mathbb{R}^{B\times L\times D}$ and $\mathbf{\Delta T}\in\mathbb{R}^{B\times L}$, where $L$ represents the maximum length of padded sequences and $B$ refers to the batch size.
3.3. State Space Model
@@
State space models are classical for modeling system dynamics, including the representative works S4, S5, and S6. SSM converts an input function $x(t)$ to an output function $y(t)$ through an implicit latent state $h(t)$ as the following steps:
(1) $h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t)$
(2) $y(t)=\mathbf{C}h(t)$
@@
where $\mathbf{A}\in\mathbb{R}^{P\times P}$, $\mathbf{B}\in\mathbb{R}^{P\times D}$, $\mathbf{C}\in\mathbb{R}^{D\times P}$ are learnable parameter matrices. To employ it on a discrete input interaction sequence $[\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}]$ instead of the continuous function $x(t)$, continuous-time SSM in Equation 1 and 2 can be discretized by a constant step size $\mathbf{\Delta}$ through zero-order hold rule:
(3) $\mathbf{h}_{t}=\bar{\mathbf{A}}\mathbf{h}_{t-1}+\bar{\mathbf{B}}\mathbf{x}_{t}$
(4) $\mathbf{y}_{t}=\bar{\mathbf{C}}\mathbf{h}_{t}$
@@
3.4. Time-Aware SSM
We introduce a time-aware SSM for explicitly capturing temporal information within sequences. It utilizes variable discretization step sizes to capture the dynamic changes of the sequence. This feature allows our method to model user-item interaction sequences in continuous time.
3.4.1. Diagonalize
Firstly, we diagonalize the system by definition $\mathbf{A}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{-1}$, to compute the linear recurrence using a parallel scan efficiently. $\mathbf{\Lambda}\in\mathbb{R}^{P\times P}$ corresponds to the diagonal matrix containing the eigenvalues and $\mathbf{V}\in\mathbb{R}^{P\times P}$ denotes the eigenvectors.
3.4.2. Discretized by Variable Timesteps
The sequence of user interactions contains various temporal intervals. These intervals establish the relative positional relationships between items, which are crucial for identifying evolving patterns of user interests. Primitive SSM implementations typically employ a constant timestep, disregarding the variable observation interval $\mathbf{\Delta T}$. Our proposed time-aware SSM discretizes the SSM with variable step $\mathbf{\Delta}$ to maintain the relative temporal relationship of the original data. It transforms the time-invariant parameters of SSM in Equation 3 and 4 into time-varying parameters:
(5) $\mathbf{h}_{t}=\bar{\mathbf{A}}_{t}\mathbf{h}_{t-1}+\bar{\mathbf{B}}_{t}\mathbf{x}_{t}$
(6) $\mathbf{y}_{t}=\bar{\mathbf{C}}\mathbf{h}_{t}$

### v2

@@ Abstract
Sequential recommendation is a key area in the field of recommender systems aiming to model user interest based on historical interaction sequences with irregular intervals. While previous recurrent neural network-based and attention-based approaches have achieved significant results, they have limitations in capturing system continuity due to the discrete characteristics. In the context of continuous-time modeling, state space model (SSM) offers a potential solution, as it can effectively capture the dynamic evolution of user interest over time. However, existing SSM-based approaches ignore the impact of irregular time intervals within historical user interactions, making it difficult to model complexed user-item transitions in sequences. To address this issue, we propose a hybrid SSM-based model called SS4Rec for continuous-time sequential recommendation. SS4Rec integrates a time-aware SSM to handle irregular time intervals and a relation-aware SSM to model contextual dependencies, enabling it to infer user interest from both temporal and sequential perspectives. In the training process, the time-aware SSM and the relation-aware SSM are discretized by variable stepsizes according to user interaction time intervals and input data, respectively. This helps capture the continuous dependency from irregular time intervals and provides time-specific personalized recommendations. Experimental studies on five benchmark datasets demonstrate the superiority and effectiveness of SS4Rec.
Index Terms:
@@
State Space Model, Continuous-Time Modeling, Sequential Recommendation.
I Introduction
@@
Recommender systems are widely employed in social media and e-commerce to alleviate information overload [1]. The core of recommender systems is accurately predicting user interests in order to suggest appropriate items for users from a large candidate pool [2]. Notably, user interests are always changing continuously over time, and this evolution is reflected in their historical interaction sequence. As a result, sequential recommender systems have been developed to offer personalized online services by capturing the users’ evolving interests through their interaction sequences.
@@
Earlier sequential recommendation approaches utilized Markov chains to capture sequential patterns by modeling the transition probabilities between items [3, 4]. With the advent of deep neural networks, a series of related methods emerged, including those based on recurrent neural networks (RNNs) [5, 6, 7], convolutional neural networks (CNNs) [8], and Transformer models [9, 10, 11]. Although these approaches achieve promising performance, they are unable to effectively capture long-range dependencies and suffer from the inference inefficiency problem. Additionally, these models are inherently discrete and designed for dealing with discrete systems, however, user interests evolve continuously and vary over time, which requires recommender systems to adapt to the user’s current state as well as the dynamic interest evolution.
@@
Recently, state space models (SSMs) have become an alternative to neural network-based models due to their effectiveness in modeling continuous dynamic systems. S4 [12] and S5 [13] utilize HiPPO matrix initialization and parallel scanning for long-range reasoning and fast inferring, respectively. The latest SSM, Mamba (i.e., S6) [14], employs an input-dependent selective state space model to retain input-relevant information while filtering out irrelevant information. This improves its sequential modeling capability and inspires researchers to explore the potential of SSMs for sequential recommendation. Mamba4Rec [15] and RecMamba [16] leverages Mamba to capture sequential dependencies from historical interactions, modeling complex relationships between users and items while partially addressing the trade-off between recommendation performance and inference efficiency.
@@
Despite the success of SSMs in efficient sequential recommendations for long sequences and large-scale datasets, these methods are still under the assumption of uniform time intervals, as Mamba was originally designed for NLP tasks. Therefore, existing SSM-based approaches focus on capturing relation-aware information but overlook the impact of varying temporal dynamics in user interest evolution. In reality, the observed interactions are often recorded as discrete time series data with irregular time intervals [17], which play a crucial role in maintaining the underlying continuity. Although some methods [18, 19, 20, 21] introduce time-dependent embeddings or kernel functions to model temporal information, they still struggle to overcome the assumption of uniform time intervals. This limitation disrupts the underlying continuity of the recommender system, impairing its ability to capture users’ evolving interests and time-specific preferences (e.g., morning vs. evening).
@@
To address these challenges, we observe that different SSM architectures possess distinct strengths and propose a hybrid SSM-based model, SS4Rec, for continuous-time sequential recommendation. SS4Rec comprises a time-aware SSM and a relation-aware SSM, each designed to capture user interests from different perspectives. The discretization steps of these two SSMs are based on the time interval sequences and the user interaction sequences, respectively. Specifically, SS4Rec first employs the time-aware SSM as an effective encoder to handle variable observation intervals, capturing time patterns. It then integrates a selective mechanism into the relation-aware SSM to extract important and relevant information, enhancing its ability to model sequential dependencies. Additionally, residual connections and layer normalization are incorporated to prevent network degradation and ensure training stability.
The major contributions of this work are as follows:
  • We formulate the sequential recommendation problem as a continuous time-varying system and address it using state space models with time-varying parameters. This work eliminates the assumption of uniform time intervals in prior work, allowing us to directly model user interest evolution with irregular-interval interactions, leading to more accurate and time-specific recommendations.
  • SS4Rec is the first work to integrate distinct parameterized state space models to simultaneously capture the continuity of recommender systems from both temporal and sequential perspectives. This design leverages the strengths of SSMs in modeling long-range dependencies while ensuring high inference efficiency.
  • Extensive experiments are conducted on five real-world datasets, and the results show the superiority and effectiveness of SS4Rec compared to other state-of-the-art models.
II Related Work
II-A Sequential Recommendation
@@
Sequential recommendation has been extensively studied, with early approaches primarily relying on Markov chain. For example, FPMC (Factorizing Personalized Markov Chains) [3] combines the factorization machine and Markov chain techniques to construct a personalized item-item relationship transition matrix for each user, enabling the next-item prediction based on users’ most recent interactions.
@@
In recent years, deep neural networks (DNNs) have gained significant attention due to their strong expressive capabilities, leading to the emergence of DNN-based models in the field of sequential recommendation. GRU4Rec [6] utilizes Gate Recurrent Unit (GRU) [22] network with ranking loss to capture the user interests within the sequence. NARM [23] applies GRU to extract the users’ general interests and intentions through a global and local encoder, respectively. SASRec [10] employs the Transformer-based architecture to model user interaction sequences, capturing long-term dependencies while adaptively adjusting attention weights of items in the sequence, thereby focusing on important information. Bert4Rec [11] introduces a Cloze task and utilizes a bidirectional attention mechanism to learn contextual information from the interaction sequences. DCRec [24] and GCL4SR [25] leverage self-supervised learning techniques to capture complex relationships within sequential data and enhance recommendation performance. All the aforementioned approaches are devoted to modeling sequential dependencies in historical user behaviors but ignore the interaction timestamps, limiting their ability to capture the dynamic evolution of user interests.
II-B Time-Aware Recommendation
@@
Several time-aware sequential recommendation approaches have been proposed to incorporate temporal information in modeling user preferences. TiSASRec [18] adopts a time-aware self-attentive model to learn the weights of different items in a sequence, integrating both absolute positions and relative time intervals between items to predict future items. TGSRec [26] combines sequential patterns and temporal collaborative signals by introducing a Temporal Collaborative Transformer (TCT) layer, capturing both user-item interactions and temporal dynamics for continuous-time recommendation. TAT4SRec [20] processes item sequences and timestamp sequences separately, using a window function to maintain continuity in timestamps. TASER [19] introduces a relative temporal attention network to capture the pairwise dependencies using absolute temporal vectors. The work of TiCoSeRec [27] demonstrates that uniformly distributed time intervals can improve performance. Different from these time-discretized approaches, we explore the continuous-time recommender systems, aiming to extract dynamic evolving information underlying continuous systems.
II-C State Space Models
@@
State space models have emerged as a promising class of architectures, demonstrating exceptional performance in sequence modeling tasks [28]. The first structured state space sequence model, S4 [12], addresses the challenge of capturing long-range dependencies through the state matrix initialization based on HiPPO framework [29]. Different from the single-input, single-output scheme of S4, S5 [13] adopts multi-input, multi-output (MIMO) SSM. By utilizing a diagonal state matrix, S5 achieves linear computational complexity and time-varying parameterization via parallel scans. The latest work of SSM, Mamba [14], incorporates a selection mechanism to filter irrelevant information while retaining relevant information, and designs a hardware-aware algorithm for both efficient training and inference. Mamba2 [30] connects attention variants with selective SSM directly for large-scale language modeling. Jamba [31] is a hybrid architecture for large-scale modeling that combines Mamba with Transformer and mixture-of-experts module to take full advantage of each architecture.
@@
Recent works have introduced Mamba into recommender systems due to its impressive performance in sequential tasks. Mamba4Rec [15] is the first to apply the Mamba block for sequential recommendation with improving performance and fast inference. RecMamba [16] replaces the Transformer block in SASRec with Mamba for lifelong sequences. MaTrRec [32] combines Mamba with Transformer, leveraging their strengths of capturing both long-range and short-range dependencies. SSD4Rec [33] introduces a bi-directional structured state-space duality block to capture contextual information, similar to Bert4Rec. Despite these advancements, the challenge of time-specific personalized recommendations with irregular intervals remains largely unexplored. By formulating this problem as a continuous time-varying system and adopting hybrid state space models to capture both temporal and sequential information of user interactions, our work sheds light on the continuous-time sequential recommendation.
III SS4Rec
III-A Framework Overview
@@
In recommender systems, the dynamic changes of user interests can be treated as a continuous time-varying system, where user-item interaction sequences depict the system from discrete states with irregular time intervals. Our proposed SS4Rec simultaneously models the relationship between user-item interactions (i.e., taking as input points of system) and captures the dynamic evolution of the continuous-time system by incorporating time-aware and relation-aware state space models (SSMs).
@@
An overview of the SS4Rec framework is shown in Fig. 1. SS4Rec is a hybrid model that combines two distinct types of SSMs for sequential recommendation. It consists of initialization, SS4Rec blocks, and continuous-time prediction. Specifically, each SS4Rec block comprises two key components: (1) Time-Aware SSM discretized by variable time steps; (2) Relation-Aware SSM featuring a selection mechanism. Additionally, residual connection and layer normalization are applied between each layer. In the following sections, we will provide a detailed description of the model architecture.
Figure 1: Overview of the SS4Rec framework.
III-B Initialization
@@
In the recommender system, $\mathbf{S}_{u}=[v_{1},v_{2},\dots,v_{n}]$ represents the item sequence of user $u$ and $\mathbf{T}_{u}=[t_{1},t_{2},\dots,t_{n}]$ represents the associated time sequence of user $u$, where $v_{i}$ denotes the ID of $i$-th item and $t_{i}$ denotes the interaction timestamp of $v_{i}$.
@@
At the beginning, the user interaction information is initialized from two aspects. First, the item sequence initialization involves mapping the sequence of items into $D$-dimensional embedding vectors using an embedding layer. In this layer, a learnable item embedding matrix $\mathbf{E}\in\mathbb{R}^{N\times D}$ is employed to project item IDs into high-dimensional vectors, where $N$ refers to the total number of items and $D$ denotes the dimension of the item embedding vectors. The embedding layer then transforms each sequence $\mathbf{S}_{u}$ into item embedding vectors $\mathbf{X}_{u}\in\mathbb{R}^{n\times D}$. The second aspect is the time interval sequence initialization. Given the time sequence $T_{u}$ of user $u$, we compute the time intervals $\Delta_{i}=|t_{i+1}-t_{i}|$ between two consecutive items $v_{i}$ and $v_{i+1}$. The time interval sequence for user $u$ is then obtained as $\mathbf{\Delta T}_{u}=[\Delta_{1},\Delta_{2},\dots,\Delta_{n}]\in\mathbb{R}^{n}$.
@@
For efficient computation, we process sequences in batches to generate the output embeddings $\mathbf{X}\in\mathbb{R}^{B\times L\times D}$ and $\mathbf{\Delta T}\in\mathbb{R}^{B\times L}$, where $L$ represents the maximum length of padded sequences and $B$ refers to the batch size.
III-C State Space Model
@@
State space models (SSMs) are classical tools for modeling system dynamics, with representative works S4, S5, and S6. A SSM converts an input function $x(t)$ to an output function $y(t)$ through an implicit latent state $h(t)$ in the following steps:
$h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t)$ (1)
$y(t)=\mathbf{C}h(t)$ (2)
@@
where $\mathbf{A}\in\mathbb{R}^{P\times P}$, $\mathbf{B}\in\mathbb{R}^{P\times D}$, $\mathbf{C}\in\mathbb{R}^{D\times P}$ are learnable parameter matrices. To employ it on a discrete input interaction sequence $[\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n}]$ instead of the continuous function $x(t)$, continuous-time SSM in (1) and (2) can be discretized by a constant step size $\mathbf{\Delta}$ through the zero-order hold rule:
$\mathbf{h}_{t}=\bar{\mathbf{A}}\mathbf{h}_{t-1}+\bar{\mathbf{B}}\mathbf{x}_{t}$ (3)
$\mathbf{y}_{t}=\bar{\mathbf{C}}\mathbf{h}_{t}$ (4)
@@
III-D Time-Aware SSM
We introduce a time-aware SSM for explicitly capturing temporal information within interaction sequences. It utilizes variable discretization step sizes to capture the dynamic changes of the sequence. This feature allows our method to model user-item interaction sequences in continuous time.
III-D1 Diagonalization
First, we diagonalize the system by definition $\mathbf{A}=\mathbf{V}\mathbf{\Lambda}\mathbf{V}^{-1}$, to compute the linear recurrence using a parallel scan efficiently. $\mathbf{\Lambda}\in\mathbb{R}^{P\times P}$ corresponds to the diagonal matrix containing the eigenvalues and $\mathbf{V}\in\mathbb{R}^{P\times P}$ denotes the eigenvectors.
III-D2 Discretization by Variable Timesteps
The sequence of user interactions contains various temporal intervals. These intervals establish the relative positional relationships between items, which are crucial for identifying evolving patterns of user interests. Primitive SSM implementations typically employ a constant timestep, disregarding the variable observation interval $\mathbf{\Delta T}$. Our proposed time-aware SSM discretizes the SSM with variable step $\mathbf{\Delta}$ to maintain the relative temporal relationship of the original data. It transforms the time-invariant parameters of SSM in (3) and (4) into time-varying parameters:
$\mathbf{h}_{t}=\bar{\mathbf{A}}_{t}\mathbf{h}_{t-1}+\bar{\mathbf{B}}_{t}\mathbf{x}_{t}$ (5)
$\mathbf{y}_{t}=\bar{\mathbf{C}}\mathbf{h}_{t}$ (6)